Select the graph of g(x)=3csc1/3x . Then describe the graph of g as a transformation of the graph of f(x)=cscx 


1
Expert's answer
2021-05-31T16:42:16-0400

Given the parent function is "f(x)=cscx"

And the final transformation function is

"g(x)=3csc\frac{1}{3}x"

The graph of parent function is:



The graph of a function "h(x)=af(x)" is the vertical stretch of the graph of (fx) if |a|>1

Let a=3

Hence the graph of a function "h(x)=af(x)=3csc(x)" is the vertical stretch of the graph of "f(x)=CSC(x)" .

The graph of function "h(x)=3csc(x)" is:



The graph function "g(x)=h(bx)" is the horizontal stretch of the graph of h(x) if 0<|b|<1

Let "b=\frac{1}{3}"

Hence the graph of a function "g(x)=h(\frac{1}{3}x)=3csc(\frac{1}{3}x)" is the horizontal stretch of the graph of "h(x)=3csc(x)" .

The graph of a function "g(x)=3csc(\frac{1}{3}x)" is